OpenStudy (anonymous):

find the angle between the given vectors to the nearest tenth of a degree. u=<8,4> , v=<9,9> options are 81.6 degrees 25.8 degrees 35.8 degrees 71.6 degrees Thankyou for the help :)

OpenStudy (anonymous):

Dot product is handy for that.

OpenStudy (anonymous):

it is multiple choice

OpenStudy (anonymous):

It sure is.

OpenStudy (anonymous):

Like CliffSedge said: \[\frac{ a * b }{ \left| a \right|\left| b \right| } = Cos \theta \] Where * = dot (dot product)

OpenStudy (anonymous):

In your case a<8,4> b<9,9> dot product is : 72 + 36 = 108 magnitude of a : \[\sqrt{8^2 + 4^2}\] magnitude of b: \[\sqrt{9^2 + 9^2}\]

OpenStudy (anonymous):

I'm not getting one of those answer choices.. How about you, @infoshark ?

OpenStudy (anonymous):

Flying out to lunch as we speak, will take a peak when I return. Sorry Cliff. Girlfriend has me well trained.

OpenStudy (anonymous):

Understood.

OpenStudy (anonymous):

@cookiemonsta , are you sure those are the correct answers for this? Did you possibly not copy down the question properly?

OpenStudy (anonymous):

I'm getting somewhere around 18 degrees, give or take, how about you @CliffSedge ?

OpenStudy (anonymous):

I don't remember, but yeah, it was something like 18.5